A tree casts a shadow of 15 meters, while a 2-meter post nearby casts a shadow of 3 meters. How tall is the tree?
Answer by checkley77(12844)

2/3:T/15
3T=2*15
3T=30
T=10 TREE IS 10M TALL.
PROOF:
2/3=10/15
30=30